101rd term of the arithmetic sequence if a1=-5 and d=-4​
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


a1 = -5 = first term
d = -4 = common difference

an = nth term of an arithmetic sequence
an = a1 + d*(n-1)
an = -5 - 4(n-1)
a101 = -5 - 4*(101-1)
a101 = -405

Answer: -405

Answer by ikleyn(52802)

.

To get the  101th term of the  AP  with  a1 = -5  and  d= -4,  you need make  100  steps
to the left in the number line,  starting from a1 =  -5;  each step length is  4  units.

It is obvious that you will be at the point   -5 - 400 = -405,  finally.

------------------

For introductory lessons on arithmetic progressions see
    - Arithmetic progressions
    - The proofs of the formulas for arithmetic progressions
    - Problems on arithmetic progressions
    - Word problems on arithmetic progressions
in this site.